Evelynmartinez14:

Jackrabbits are capable of reaching speeds up to 40 miles per hour. How fast is this in feet per second? (Round to the nearest whole number.) 5,280 feet = 1 mile

dude:

They already gave you the conversion, you can use fractions to set this up You want to multiply this number by what was given above (5,280 feet = 1 mile) Then convert hours to seconds using standard conversion measurements [60 minutes in an hour --- 60 seconds in a minute] \(\dfrac{211,200miles}{\color{red}{\cancel{1hr}}}\times \dfrac{\color{red}{\cancel{1hr}}}{60min}=\dfrac{\#mi}{1min}\) [# is what you get from dividing 211,200 by 60] \(\dfrac{\#mi}{1min}\times \dfrac{1min}{60 seconds}= mi/seconds\) [Divide by 60 once again with the number you got from the first part. This should be your answer]